a pipe of radius 7m contain natural oil of density 800 and moves at velocity 25m/s the pipe is branched into 2 tubes of A and B if ra equal 2m and rb equal 4m also va equal 10m/s if there was a leakage in the pipe of flow 2m³/s what's the mass of flowing in the pipe in 10 min and what is the velocity of the tube B

1
Expert's answer
2021-11-24T19:14:04-0500
"\pi (7)^2(25)=\pi (2)^2(10)+\pi (4)^2v_b+2\\v_b=74\frac{m}{s}"

"m=\pi (7)^2(25)(800)(600)=1.85\cdot10^9\ kg"